# Probabilistic Robustness Analysis in High Dimensional Space: Application to Semantic Segmentation Network

### v1

@@ 2 Preliminaries
3 Scaling Probabilistic Reachability Analysis on SSN with Strong Guarantees
Given a specific distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ for sampling inputs $x$ from the input set $\mathbf{I}$, reasoning about the resulting output distribution $y\sim\mathcal{Y}$, is infeasible due to the high degree of nonlinearity in neural networks. As a result, providing probabilistic coverage guarantees over the network’s output space becomes a significant challenge. A key advantage of conformal inference is its distributional robustness—it produces valid guarantees that are robust to all family of distributions that can well represent the samples collected in the calibration dataset. This property makes conformal inference particularly well-suited for reachability analysis of neural networks, which are known for producing complex and often intractable output distributions.
However, a major challenge arises when applied to high-dimensional datasets: The scalar elements $R$ in calibration dataset are defined over the outputs $y\sim\mathcal{Y}$ and the space of all possible distributions $\mathcal{Y}$ that can adequately represent the calibration dataset is significantly larger in higher dimensions and thus conformal inference tends to be overly conservative. This poses a significant obstacle when applying CI to semantic segmentation networks (SSNs), which are known for producing extremely high-dimensional outputs. In this paper, we propose a novel learning-based neural network reachability analysis method for SSNs, that leverages deflative principal component analysis (PCA) (Mackey, 2008) to address the dimensionality challenges inherent in conformal inference when applied to provide coverage guarantees on neural network with high dimensional outputs.
In this section, we present our reachability technique in an incremental manner to clearly illustrate both the contributions and the challenges involved in applying conformal inference to SSNs. We begin by introducing a naive baseline approach, then highlight its limitations on high-dimensional spaces. Subsequently, we describe the enhancements we propose to overcome these challenges, culminating in the final version of our reachability algorithm.
3.1 Naive Reachability Technique via Conformal Inference
The first step to find a probabilistic reachset on neural networks with conformal inference (CI) is to generate a calibration dataset $\mathbf{M}$ of size $m$ and a training dataset $\mathbf{T}$ of size $t$.
Generating train dataset. We sample $t$ different inputs $x_{j}^{\text{train}},j=1,2,\ldots,t$ from $\mathbf{I}$, with any distribution of interest, $x\stackrel{{\scriptstyle\mathcal{W^{\prime}}}}{{\sim}}\mathbf{I}$, and compute their corresponding outputs $\mathbf{vec}(y_{j}^{\text{train}})=f(\mathbf{vec}(x_{j}^{\text{train}}))$. This gives us the training dataset, $\mathbf{T}=\left\{(x_{1}^{\text{train}},y_{1}^{\text{train}}),(x_{2}^{\text{train}},y_{2}^{\text{train}}),\ldots,(x_{t}^{\text{train}},y_{t}^{\text{train}})\right\}$.
Generating Calibration dataset. We sample the input dataset with $m$ different inputs $x^{\text{calib}}_{i},i=1,2,\ldots,m$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and their corresponding set of outputs $\mathbf{vec}(y^{\text{calib}}_{i})=f(\mathbf{vec}(x^{\text{calib}}_{i})),i=1,2,\ldots,m$. This gives us the input/output dataset $\mathbf{IO}=\left\{(x_{1}^{\text{calib}},y^{\text{calib}}_{1}),(x_{2}^{\text{calib}},y^{\text{calib}}_{2}),\ldots,(x_{m}^{\text{calib}},y^{\text{calib}}_{m})\right\}$. We next attempt to use $\mathbf{IO}$ to compute a suitable nonconformity score, $R\in\mathbb{R}_{\geq 0}$. To that end, inspired by recent work on using conformal inference in time-series Cleaveland et al. (2024), we design the nonconformity scores as follows in Equation (7). This design provides us a hyper-rectanglular set. Assuming the output $\mathbf{vec}(y)=\left[\mathbf{vec}(y)(1),\ldots,\mathbf{vec}(y)(n)\right]\in\mathbb{R}^{n}$, we propose, the nonconformity scores $R^{\text{calib}}_{i}$ as,
$R^{\text{calib}}_{i}=\max\left(\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{calib}}_{i})(n)-c(n)|}{\tau_{n}}\right),\quad i=1,2,\ldots,m$ (7)
where $c\in\mathbb{R}^{n}$ is the average of the members in the train dataset $\mathbf{T}$, and for $k=1,2,\ldots,n$, the normalization factor $\tau_{k}$ is proposed to normalize the elements of the random vectors $|\mathbf{vec}(y)-c|$ which can also be obtained using the train dataset as,
$\tau_{k}:=\max\left(\ \ \tau^{*},\max\left(\ \ \ |\mathbf{vec}(y_{1}^{\text{train}})(k)-c(k)|,\ \ \ldots,\ \ |\mathbf{vec}(y_{t}^{\text{train}})(k)-c(k)|\ \ \ \right)\ \ \right),$ (8)
where we set $\tau^{*}=10^{-5}\left(\sum_{k=1}^{n}\sum_{j=1}^{t}|\mathbf{vec}(y_{j}^{\text{train}})(k)-c(k)|\right)/nt$, to avoid facing zero as our normalization factor. We then propose the calibration dataset as $\mathbf{M}=\left\{R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}\right\}$.
The next step is to sort the nonconformity scores in $\mathbf{M}$ based on their magnitude. Without loss of generality let’s assume $R^{\text{calib}}_{1}<R^{\text{calib}}_{2}<\ldots<R^{\text{calib}}_{m}$. Given a new draw $x^{\text{unseen}}$ from the distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ and the architecture of the neural network, the output $\mathbf{vec}(y^{\text{unseen}})=f(\mathbf{vec}(x^{\text{unseen}}))$ follows a specific distribution. We call this distribution as $y\sim\mathcal{Y}$. Furthermore given the map we proposed to compute the nonconformity scores, the i.i.d random variables,
$R^{\text{unseen}}=\max\left(\frac{|\mathbf{vec}(y^{\text{unseen}})(1)-c(1)|}{\tau_{1}},\ldots,\frac{|\mathbf{vec}(y^{\text{unseen}})(n)-c(n)|}{\tau_{n}}\right)$ (9)
adopt another distribution where we call it $R\sim\mathcal{D}$. Here $R^{\text{unseen}}$ and the elements $R^{\text{calib}}_{1},R^{\text{calib}}_{2},\ldots,R^{\text{calib}}_{m}$ are all sampled from the distribution $\mathcal{D}$. This means that we can apply conformal inference. Therefore, for the new draw $R^{\text{unseen}}$, given a desired rank, $\ell\leq m$ and a desired miscoverage level $\epsilon$, we can propose the following $\langle\epsilon,\ell,m\rangle$ guarantee,
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R^{\text{calib}}_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$
Following the definition in Equation (9), the logical statement $P_{1}(\ell,m):=[R^{\text{unseen}}\leq R_{\ell}^{\text{calib}}]$ is logically equivalent to
$P_{2}(\ell,m):=\bigwedge_{k=1}^{n}\left[c(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\sigma_{k}\right],\quad\text{where}\ \sigma_{k}=\tau_{k}R^{\text{calib}}_{\ell},$ (10)
@@ 3.1 Naive Reachability Technique via Conformal Inference
which defines a hyper-rectangular set as the reachset for any unseen output of the neural network generated by $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\ \mathbf{I}$. Since $P_{1}(\ell,m)$ satisfies the $\langle\epsilon,\ell,m\rangle$ guarantee, this reachset also provides the same coverage guarantee on the unseen outputs of the neural network. Appendix C provides a clarifying example on this technique demonstrating that it is effective even on very deep neural networks, offering strong probabilistic guarantees with a reasonable number of required sampling. As the architecture of a deep neural network becomes more complex, the distribution of its output can exhibit increasingly intricate behavior. Conformal inference remains robust to any output distribution that represents well the calibration dataset, allowing it to handle such complexities reliably. However, the tightness of the reachset does not persist when the dimensionality of the network’s output increases.
@@
  • The second issue is that the space of distributions $\mathcal{Y}$ that can provide some distribution $\mathcal{D}$ which well represent the calibration dataset is significantly larger in high dimensional space.
3.2 From Hyper-Rectangular Reachsets to Convex Hull Representations
Our approach to constructing a general convex hull as the reachset is based on approximating the original deep neural network $f$ with another computation graph $g:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, for which a guaranteed deterministic reachable set can be derived. The model $g$ is generated based on the training dataset $\mathbf{T}$, which implies it is accurate only for inputs $x$ sampled from the set $\mathbf{I}$. The core idea is to utilize its deterministic reachable set, and then inflate it to account for approximation error between $f$ and its surrogate $g$, thereby obtaining a probabilistic reachset for the original model $f$ with an $\langle\epsilon,\ell,m\rangle$ guarantee.
Obtaining the surrogate model $g$ serves a dual purpose: it not only provides an approximation for $f$, but also defines the shape of a convex hull which when inflated serves as a tighter over-approximation of the probabilistc reachset of $f$. Once $g$ is computed, this convex shape can be computed as a convex hull. To inflate this surrogate reachset, we follow the procedure outlined in the naive technique (Section 3.1) to generate a hyper-rectangle that bounds the prediction error with a provable $\langle\epsilon,\ell,m\rangle$ guarantee. In other words, we define $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$ and apply the naive technique on $q(\mathbf{vec}(x))$ to compute its hyper-rectangular reachset with $\langle\epsilon,\ell,m\rangle$ guarantee. We refer to this hyper-rectangle as inflating set.
In conclusion, if we denote the deterministic reachset of surrogate model as the starset $\mathcal{R}_{g}(\mathbf{I})$,
@@ 3.2 From Hyper-Rectangular Reachsets to Convex Hull Representations
This technique offers two key improvements over the naive approach presented in Section 3.1:
@@
To obtain a surrogate model $g$ with high fidelity while ensuring a deterministic reachable set, we propose using the convex hull $\mathbf{CH}$ over the training outputs $y_{j}^{\text{train}},\;j=1,\dots,t$,
$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
and then incorporating a clipping block (see Appendix D) that projects the outputs of $f$ for any unseen image $x^{\text{unseen}}$, into this convex hull. Denoting the clipping block with $\mathbf{CLP}$ we propose the surrogate model as: $g(\mathbf{vec}(x))=\mathbf{CLP}(f(\mathbf{vec}(x)))$, that makes certain $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$.
Despite the improvements introduced in this section, reachability analysis remains unscalable in high-dimensional settings. The fundamental difficulty lies in the fact that projecting $\mathbf{vec}(y^{\text{unseen}})$ on $\mathbf{CH}$ does not scale well with the dimensionality of the output space. To mitigate this, we incorporate an scalable version of Principal Component Analysis (PCA), known as defelation algorithm, before clipping, a technique known to be effective for handling PCA in high-dimensional spaces. Therefore, the error $q(x)$ primarily arises from the inclusion of PCA and clipping block.
@@
Initialize $z_{1}\ ,\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ \mathbf{vec}(y_{1}^{\text{train}}),\ \mathbf{vec}(y_{2}^{\text{train}}),\ \ldots,\ \mathbf{vec}(y_{t}^{\text{train}})$$A\leftarrow\ \left[\ \right]$
@@ 3.3 Surrogate Model for SSNs via Principal Component Analyis
We utilize this matrix to represent the high dimensional logits $y_{j}^{\text{train}}$ with a lower dimentional representative $v_{j}\in\mathbb{R}^{N}$ where $v_{j}=A^{\top}\mathbf{vec}(y_{j}^{\text{train}})$. Principal Component Analysis is a technique to reduce the dimensionalty, but it does not scale to high dimensional spaces. To address this issue, deflation algorithms have been proposed (see Algorithm 1) that compute principal components iteratively, in order of significance (Mackey, 2008). At each iteration, once the most dominant principal component is identified, the dataset is projected onto the orthogonal complement of that direction—effectively removing its contribution—before computing the next principal direction.
In this paper, we adopt a learning based deflation algorithm that is presented in Algorithm 1. In this algorithm we set the components of the principal direction as trainable parameters and maximize the variance of the dataset in the direction of this vector. At each iteration of this algorithm, the optimization to train the principal direction have one global maxima one global minima and n2n-2 saddle points, which $n$ is the dimension of logits $\mathbf{vec}(y_{j}^{\text{train}})$ Mackey (2008). Therefore, although the training operates in a high-dimensional space, convergence to global maxima is guaranteed.
@@
Detecting Robust, Nonrobust and Unknown Pixels with Strong Probabilistic Guarantees
Once the reachset over the SSN logits $y\in\mathbb{R}^{n}$ is constructed as a starset, we project it onto each logit component. This allows us to determine the range of values for each class label $l\in\mathbf{L}$ at every pixel location $(i,j)\in\{1,\ldots,h\}\times\{1,\ldots,w\}$. Consequently, each pixel $(i,j)$ is associated with $L$ logit intervals, one for each class. Let $l^{*}\in\mathbf{L}$ denote the class whose logit interval has the highest lower bound at pixel $(i,j)$. If this lower bound is not strictly greater than the upper bounds of all other classes’ intervals, the pixel is labeled as unknown as it can obtain more than one labels under UBAA. In case it is greater, then if $l^{*}=\mathsf{SSN}(x)(i,j)$, the prediction at that pixel is labeled as robust and if $l^{*}\neq\mathsf{SSN}(x)(i,j)$, it is labeled as non-robust. The Algorithm 2 explains this procedure in detail.
4 Experiments
In our numerical evaluation, we pose three different research questions and address each of them using numerical results. The main one is available in this section and the rest of them are in Appendix E. We utilized a Linux machine, with $48$ GB of GPU memory, $512$ GB of RAM, and $112$ CPUs.
@@ 4 Experiments
// $\mathtt{Project\ the\ Starset}$

### v2

@@ 2 Preliminaries
3 Scaling Probabilistic Reachability Analysis on SSN with Strong Guarantees Using Clipping-Block
@@
As the architecture of a deep neural network becomes more complex, the distribution of its output can exhibit increasingly intricate behavior. Appendix B summarizes the naive technique proposed by hashemiscaling for the reachability of SSNs. As it is explained in Appendix B, this technique for reachability via conformal inference remains robust to any output distribution that represents well a sampled calibration dataset, allowing it to handle such complexities reliably. However, the tightness of this reachset does not persist when the dimensionality of the network’s output increases.
@@
  • The second issue is that the space of distributions $\mathcal{Y}$ that can provide some distribution $\mathcal{D}$ which well represent a sampled calibration dataset is significantly larger in high dimensional space, resulting in a conservative reachable set.
To tackle these challenges, hashemiscaling propose constructing a more flexible reachset in the form of a general convex hull, rather than a simple hyper-rectangle. This is achieved by approximating the original neural network $f$ with an alternative computation graph $g:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n}$, for which a guaranteed deterministic reachable set can be derived. To obtain the surrogate model $g$ they suggest training a small ReLU neural network over the train dataset $\mathbf{T}$ and then compute its deterministic reachset over $\mathbf{I}$ using NNV toolbox tran2020nnv. Since $g$ is is trained on $\mathbf{T}$, we infere it is accurate only for inputs $x$ sampled from the set $\mathbf{I}$. The core idea is to utilize its deterministic reachable set (also known as surrogate reachset), and then inflate it to account for approximation error between $f$ and its surrogate $g$, thereby obtaining a probabilistic reachset for the original model $f$ with an $\langle\epsilon,\ell,m\rangle$ guarantee.
To inflate the surrogate reachset, hashemiscaling adopt the procedure from the naive approach to construct a hyper-rectangle that bounds the prediction error while maintaining a provable $\langle\epsilon,\ell,m\rangle$ guarantee. Concretely, let $q(\mathbf{vec}(x))=f(\mathbf{vec}(x))-g(\mathbf{vec}(x))$. The naive technique is then applied to $q(\mathbf{vec}(x))$ to compute its hyper-rectangular reachset under the same $\langle\epsilon,\ell,m\rangle$ guarantee. This resulting hyper-rectangle is referred to as the inflating set. In conclusion, if we denote the deterministic reachset of surrogate model as $\mathcal{R}_{g}(\mathbf{I})$,
@@
This technique offers two key improvements over the naive approach:
@@
However, the small ReLU neural network proposed in hashemiscaling introduces three main challenges for this methodology:
  • The surrogate model may suffer from fidelity issues, yielding correct yet conservative probabilistic reachable sets.
  • As the perturbation dimensionality grows, the ReLU surrogate’s input dimension increases, creating scalability issues for training. In addition, deterministic reachability on the ReLU surrogate also scales poorly, often leading to out-of-memory errors in the NNV toolboxtran2020nnv. Finally, because NNV is restricted to polytopes, image perturbations must be constrained to $\ell_{\infty}$.
In the next section, we introduce the idea of clipping-block for constructing the surrogate model $g$. Unlike the previous method, our new approach requires no additional training—eliminating fidelity concerns—and imposes no restrictions on the perturbation of the input image.
3.1 Generating the Surrogate Model via Clipping-Block
$\mathbf{CH}=\{z\mid z=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\ \ \sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0\}$
that for $l=2$ is a quadratic programming and for $l=1,\infty$ is a linear programming. Among the three options, we are most interested in $l=\infty$, since it minimizes the maximum error across the logit components, making it a suitable choice for reachability analysis. In addition, the $l=2$ formulation is far more memory-intensive compared to the $l=\infty$ case, and LP formulation scales more gracefully. Each problem is linear, and batching helps amortize model builds, and parallelism is also effective — which is why we observe it running stably on our setup.
Thus, our surrogate model $g$ is composed of two main components. The first is the base model $f$, which takes the vectorized image $\mathbf{vec}(x^{\text{unseen}})$ and produces the corresponding logit vector $\mathbf{vec}(y^{\text{unseen}})$. The second is the clipping block, which takes $\mathbf{vec}(y^{\text{unseen}})$ as input and returns its projection onto the convex hull $\mathbf{CH}$, denoted as $\mathbf{vec}(\hat{y}^{\text{unseen}})$. In other words, if we denote the projection operation performed in the clipping block with $\mathbf{CLP}$, then,
$g(\mathbf{vec}(x^{\text{unseen}}))=\mathbf{vec}(\hat{y}^{\text{unseen}})=\mathbf{CLP}(f(\mathbf{vec}(x^{\text{unseen}})))$
and obviously $\mathbf{CH}$ represents the deterministic reachable set of the surrogate model, $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}$. The availability of the surrogate reachset, regardless of the type or dimensionality of the input perturbation, removes the need to assume sparsity and eliminates any restrictions on the form of the image perturbation. Appendix C uses a toy example to make a comparison between our clipping block approach and the Naive technique on the reachability of deep neural networks.
To generate the convex hull, it is not necessary to include the entire training dataset. Instead, one can identify the extreme points sartipizadeh2016computing and construct the convex hull using only those points, which reduces the number of coefficients $\alpha$ without sacrificing accuracy. This approach makes the convex hull more efficient to be used as it reduces the number of coefficients $\alpha$. Existing algorithms for identifying extreme points include Quickhull barber1996quickhull and Clarkson’s algorithm clarkson1988applications. Their computational complexities are $\mathcal{O}(t\log(t))$ for $n=2,3$ and $\mathcal{O}(t^{\lfloor n/2\rfloor})$ for $n\geq 4$, which do not scale well to high-dimensional spaces. Another approach to this problem is to approximate the convex hull by allowing a controlled level of inaccuracy sartipizadeh2016computing; jia2022investigating. However, as reported in the literature casadio2025nlp, this method has not shown promising performance even in high-dimensional settings such as large language models, whose dimensionality is still lower than that of semantic segmentation networks.
For this reason, we keep the original formulation and retain all points in the training dataset when constructing the convex hull. Importantly, this does not create scalability issues for our experiments, since conformal inference is highly data-efficient. For example, guarantees of order $\delta_{1}=0.999$ and $\delta_{2}=0.997$ can be verified with a calibration dataset of size $m=8000$. Because we consider the size of the training dataset to be at most $50\%$ of the calibration dataset size, then $t=4000$ training samples suffice for our experiments, requiring only $4000$ coefficients $\alpha$ in the convex hull formulation. Moreover, since the number of extreme points grows exponentially with the dimension of the logits, i.e., $\mathcal{O}((\log t)^{n-1})$ dwyer1988average, it is highly likely that most training samples will serve as extreme points. This renders attempts to distinguish between extreme and interior points largely ineffective.
Similar to hashemiscaling, which encounters scalability issues when training a surrogate model in high-dimensional output spaces, our approach also suffers from this limitation. Specifically, projecting $\mathbf{vec}(y^{\text{unseen}})$ onto $\mathbf{CH}$ using linear programming does not scale efficiently with output dimensionality. To address this, the authors of hashemiscaling introduce a scalable variant of Principal Component Analysis (PCA), known as the deflation algorithm, prior to training. We also found this strategy beneficial and adopt PCA before implementing the clipping block. Consequently, we conclude that the prediction error in our method arises primarily from the combined effects of PCA and the clipping block. The following section details this incorporation, which enables our technique to extend to robustness analysis of SSNs.
@@ 3.1 Generating the Surrogate Model via Clipping-Block
Initialize $z_{1}\ ,\ z_{2},\ \ldots,\ z_{t}\ \leftarrow\ \mathbf{vec}(y_{1}^{\text{train}}),\ \mathbf{vec}(y_{2}^{\text{train}}),\ \ldots,\ \mathbf{vec}(y_{t}^{\text{train}})$ , $A\leftarrow\ \left[\ \right]$
@@ 3.2 Surrogate Model for SSNs via Principal Component Analyis
We employ this matrix to compress the high-dimensional logits $y_{j}^{\text{train}}$ into a reduced representation $v_{j}\in\mathbb{R}^{N}$, defined as $v_{j}=A^{\top}\mathbf{vec}(y_{j}^{\text{train}})$. While Principal Component Analysis (PCA) is the standard tool for dimensionality reduction, its direct application does not scale effectively in very high-dimensional settings. To overcome this limitation, deflation-based methods have been developed (Algorithm 1). These algorithms extract principal components one at a time, ordered by significance (mackey2008deflation). After each dominant component is obtained, the dataset is projected onto its orthogonal complement, thereby eliminating its influence before proceeding to the next iteration.
In this paper, we employ a learning-based deflation algorithm, outlined in Algorithm 1. In this approach, the components of the principal direction are treated as trainable parameters, and the objective is to maximize the variance of the dataset along this direction. At each iteration, the optimization problem for learning the principal direction admits one global maximum, one global minimum, and n2n-2 saddle points, where $n$ denotes the dimension of the logits $\mathbf{vec}(y_{j}^{\text{train}})$ mackey2008deflation. Consequently, despite operating in a high-dimensional space, convergence to the global maximum is guaranteed.
@@
$\mathbf{HR}=\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\{z\mid z=c+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\qquad\sum_{i=1}^{n}\beta_{i}=1,\quad\beta_{1},\cdots,\beta_{n}>0\}$
$\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\mathbf{CH}\oplus\mathbf{HR}=$
$\{z\mid z=c+\sum_{j=1}^{t}\alpha_{j}A\ v^{\text{train}}_{j}+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\ \ \alpha_{1},\cdots,\alpha_{t},\beta_{1},\cdots,\beta_{n}\geq 0,\ \sum_{j=1}^{t}\alpha_{j}=1,\ \sum_{i=1}^{n}\beta_{i}=1\}$
The projection of the probabilistic reach set $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$ onto each class of a given pixel can be performed efficiently. To this end, we collect the training points $v_{j}^{\text{train}},;j=1,\ldots,t$. Since the mapping from latent space to the original space is linear ($\mathbf{vec}(\hat{y})=Av$), the vertices of the convex hull in latent space are preserved under this transformation. In other words, if $v^{*}$ is an extreme point of the convex hull in latent space, then $Av^{*}$ is an extreme point of the convex hull in the original space. Therefore, we collect the points $\mathbf{vec}(\hat{y}_{j}^{\text{train}})=Av_{j}^{\text{train}}$ and compare them to determine the upper bound $\mathbf{ub}$ and lower bound $\mathbf{lb}$ of the surrogate model’s predictions. Consequently, the clipping block in the surrogate model ensures that every unseen point $\mathbf{vec}(\hat{y}^{\text{unseen}})$ satisfies, $\mathbf{vec}(\hat{y}^{\text{unseen}})\in[\mathbf{lb},\mathbf{ub}]$. For a given pixel and class, let the corresponding logit value be the $k$-th component of $\mathbf{vec}(y^{\text{unseen}})$. Then the projection of the reachset on this component satisfies: $\mathbf{vec}(y^{\text{unseen}})(k)\in[c(k)+\mathbf{lb}(k)-\sigma_{k}\ \ ,\ \ c(k)+\mathbf{ub}(k)+\sigma_{k}]$.
We stress that setting $\mathcal{W}^{\prime}=\mathcal{W}$ and $t=m/2$ markedly improves the surrogate model’s fidelity in our new approach a setting we always follow in our experiments via clipping block technique.
@@
1mm // $\mathtt{Project\ the\ reachset}$
@@
4 Experiments
In our numerical evaluation, we pose two different research questions and address each of them using numerical results. The main one is available in this section and the rest of them are in Appendix D. We utilized a Linux machine, with $48$ GB of GPU memory, $512$ GB of RAM, and $112$ CPUs.